OpenStudy (anonymous):

Solve. A. –3 B. 3 C. –27 D. 27

OpenStudy (anonymous):

-1/3 <-- thats a fraction b = 9

OpenStudy (anonymous):

@abigailsimmons

OpenStudy (anonymous):

@camrondoherty

OpenStudy (jordanloveangel):

one sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i know how to do it i just need to know how can you tell if its a - or the + one

OpenStudy (anonymous):

but my answer was -27

OpenStudy (anonymous):

-1/3 b = 9

OpenStudy (jordanloveangel):

yes i got the same thing

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

thank you

OpenStudy (jordanloveangel):

yeah np u still need help

OpenStudy (anonymous):

well idk i'm still tryna do one if i need help i'll tell ya

OpenStudy (anonymous):

:)

OpenStudy (jordanloveangel):

oh kk

OpenStudy (anonymous):

Solve. –30 = 6z A. –5 B. –6 C. –24 D. –180

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

why does it do that

OpenStudy (anonymous):

i picked -5

OpenStudy (anonymous):

by the way next to the answers there a - to all of them

OpenStudy (anonymous):

jordan?

OpenStudy (jordanloveangel):

yes sorry one sec

OpenStudy (anonymous):

its okay

OpenStudy (jordanloveangel):

Simplifying -30 = 6z Solving -30 = 6z Solving for variable 'z'. Move all terms containing z to the left, all other terms to the right. Add '-6z' to each side of the equation. -30 + -6z = 6z + -6z Combine like terms: 6z + -6z = 0 -30 + -6z = 0 Add '30' to each side of the equation. -30 + 30 + -6z = 0 + 30 Combine like terms: -30 + 30 = 0 0 + -6z = 0 + 30 -6z = 0 + 30 Combine like terms: 0 + 30 = 30 -6z = 30 Divide each side by '-6'. z = -5 Simplifying z = -5

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thats what i got

OpenStudy (jordanloveangel):

cool

OpenStudy (anonymous):

What is the value of n? 40 = - 5n/8

OpenStudy (anonymous):

n =____

OpenStudy (anonymous):

-64?

OpenStudy (jordanloveangel):

one sec

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

40=-5n/8 Simple and best practice solution for 40=-5n/8 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for 40=-5n/8 equation: n in (-oo:+oo) 40 = (-5*n)/8 // - (-5*n)/8 40-((-5*n)/8) = 0 (5/8)*n+40 = 0 5/8*n+40 = 0 // - 40 5/8*n = -40 // : 5/8 n = -40/5/8 n = -64 n = -64

OpenStudy (anonymous):

:) thanx

OpenStudy (anonymous):

k last one

OpenStudy (jordanloveangel):

np

OpenStudy (jordanloveangel):

please post it

OpenStudy (anonymous):

and okay :)

OpenStudy (anonymous):

you want me to close this one

OpenStudy (anonymous):

and say it in another one

OpenStudy (jordanloveangel):

yes

OpenStudy (anonymous):

k

OpenStudy (jordanloveangel):

thanx